# spacing=nonfrench

### v2

@@ Appendix A Proof of $\epsilon$-LDP with the Laplacian mechanism and Post-Processing Property
Appendix B Hyperparameter choices and dataset preparation
In this section, we document hyperparameter choices and dataset splits for the experiments in Section 4. We use three image benchmarks for evaluation: SVHN with 10 classes of 32x32x3 cropped street view house number images, Fashion-MNIST with 10 classes of 28x28 greyscale fashionable clothing images, and PathMNIST with 9 classes of 28x28x3 medical images of pathology. We train ${\mathcal{M}_{\rm R}}$ with a learning rate of 0.1 for 200 epochs on Fashion-MNIST and SVHN and 10 epochs on PathMNIST. ${\mathcal{M}_{\rm R}}$ is trained on 35k data points for Fashion-MNIST, 48,257 data points for SVHN and 89,996 data points for PathMNIST. ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ are split from the remaining data points unseen by ${\mathcal{M}_{\rm R}}$, where ${\mathcal{D}_{\rm priv}}$ is used to train and evaluate ${\mathcal{M}_{\rm L}}$, and ${\mathcal{D}_{\rm val}}$ is used for pure model generalizability evaluation. Specifically, CIFAR-10 has 15k data points in the candidate pool of ${\mathcal{D}_{\rm priv}}$, Fashion-MNIST and SVHN have 25k, and PathMNIST has 10,004. As for ${\mathcal{D}_{\rm val}}$, its size is 10k for both CIFAR-10 and Fashion-MNIST, 26,032 for SVHN, and 7,180 for PathMNIST. ${\mathcal{D}_{\rm infer}}$ in Sections 4.2 and 4.5 has a size of $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$ (around 2.2M). In Section 4.4, $|{\mathcal{D}_{\rm priv}}|$ is set to 125, 250, 500, 1k and 1.5k, whereas $|{\mathcal{D}_{\rm infer}}|$ is set to 15,500, 62,250, 250k and 500k, respectively. In this section, another set of comparison experiment is done with $|{\mathcal{D}_{\rm priv}}|$ fixed at 1.5k and ${\mathcal{D}_{\rm infer}}$ has a varying size of 250k, 500k, 1M and 2M (i.e., all possible superimposition pairs like Section 4.2). ${\mathcal{M}_{\rm L}}$’s learning rate is 0.1 for Fashion-MNIST and 0.001 for SVHN and PathMNIST. The number of training epochs is set to 15 across all the datasets in all scenarios. For the latent space analysis, we train VAE for 200 epochs with a learning rate set to 0.001 for Fashion-MNIST. The number of epochs is set to 300 and the learning rate is 0.0001 for SVHN and PathMNIST.
Appendix C Examples of ${\mathcal{D}_{\rm infer}}$ data with LDP noise
Figures 11, 12, and 13 present comparisons between the original and noised data samples under different levels of $\epsilon$-LDP noise. They also illustrate the impact of our two noise application mechanisms, LDPKiT-Rand and LDPKiT-Sup.
Appendix D Additional experimental results
This section is complementary to the evaluation results presented in Section 4.
D.1 Final accuracy on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$
@@
Tables 5, 6 and 7 show the tabulated final accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ of SVHN, Fashion-MNIST and SVHN that SIDP, LDPKiT-Rand and LDPKiT-Sup can achieve, which are also presented as bar graphs in Figures 4 and 5 in Section 4.2 and Figures 10, 9, 15 and 14 in Section 4.5 and Appendix D.2. We record the accuracies at the last epoch of training. The results draw the same conclusion as Section 4 that LDPKiT helps offset the accuracy trade-offs brought by LDP noise. The tables also show that LDPKiT can almost always achieve higher inference accuracy than SIDP, but the advantage is less obvious when the least noise is added (i.e., $\epsilon=30$ or $\epsilon=15$). However, there is barely privacy protection in these cases, as SIDP accuracies are also very high, indicating a significant information leakage to ${\mathcal{M}_{\rm R}}$. As more LDP noise is added to ${\mathcal{D}_{\rm priv}}$ to preserve privacy, the gap in utility that LDPKiT provides over SIDP also increases. Therefore, LDPKiT offers greater benefits in regimes with stronger privacy protection and correspondingly more noise.
D.2 Impact of $\epsilon$ on MobileNetV2’s Accuracies on ${\mathcal{D}_{\rm priv}}$ versus ${\mathcal{D}_{\rm val}}$
We observe the same trend in Figures 14 and 15 as per discussion in Section 4.5 with MobileNetV2 across all three benchmarks.
D.3 Additional Dataset Size Sensitivity Analysis
@@
The results in this section are complementary to the discussion in Section 4.4. Table 8 shows the impact of $|{\mathcal{D}_{\rm infer}}|$ and $|{\mathcal{D}_{\rm priv}}|$ with other $\epsilon$ values. All the additional results align with the conclusions drawn from Table 3. Figure 16 demonstrates the impact of $|{\mathcal{D}_{\rm infer}}|$ with other $\epsilon$ values on the three image benchmarks. The takeaway aligns with the conclusion drawn from Figure 8 that although larger ${\mathcal{D}_{\rm infer}}$ helps model training and leads to higher accuracies, LDPKiT can already achieve satisfactory accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ when $|{\mathcal{D}_{\rm infer}}|$ is 500k. The main takeaway aligns with the discussion in Section 4. LDPKiT’s accuracies increase as $|{\mathcal{D}_{\rm infer}}|$ increases; however, the improvement becomes minor with larger ${\mathcal{D}_{\rm infer}}$ and the accuracies are sufficiently high when $|{\mathcal{D}_{\rm infer}}|$ is 500k. Furthermore, LDPKiT is less sensitive to $|{\mathcal{D}_{\rm priv}}|$, so LDPKiT can achieve decent accuracies when the user owns a relatively smaller ${\mathcal{D}_{\rm priv}}$.
@@
Table 9: Euclidean distances between centroids of clusters shown in Figure 17 on PathMNIST.
@@
17(a) LDPKiT-Sup (4,4) 1.9453 1.5981 2.5698
LDPKiT-Rand 4 1.8724 2.5812 3.3226
17(b) LDPKiT-Sup (8,4) 1.4984 2.1860 2.0151
LDPKiT-Rand 8 1.6577 3.3427 3.0580
17(c) LDPKiT-Sup (8,8) 1.6301 2.4562 1.5323
LDPKiT-Rand 8 1.6407 3.3359 3.0413
@@
Table 10: Euclidean distances between centroids of clusters shown in Figure 18 on SVHN.
@@
18(a) LDPKiT-Sup (1,1) 2.6877 3.1031 3.2987
LDPKiT-Rand 1 4.4212 3.0084 3.1381
18(b) LDPKiT-Sup (1,8) 2.8854 3.0853 2.9836
LDPKiT-Rand 1 4.4242 3.0094 3.1386
18(c) LDPKiT-Sup (2,8) 3.8513 2.5689 3.1065
LDPKiT-Rand 2 5.2321 3.0173 3.5754
@@
Table 11: Normalized Zest distance results with $l_{1}$ distance metric on ${\mathcal{M}_{\rm R}}$ and ${\mathcal{M}_{\rm L}}$.
@@
$\epsilon$=2.0 $\epsilon$=1.5 $\epsilon$=1.5 $\epsilon$=1.25 $\epsilon$=10.0 $\epsilon$=7.0
@@
Table 12: Normalized Zest distance results with $l_{2}$ distance metric on ${\mathcal{M}_{\rm R}}$ and ${\mathcal{M}_{\rm L}}$.
@@
$\epsilon$=2.0 $\epsilon$=1.5 $\epsilon$=1.5 $\epsilon$=1.25 $\epsilon$=10.0 $\epsilon$=7.0
@@
Table 13: Normalized Zest distance results with $l_{\infty}$ distance metric on ${\mathcal{M}_{\rm R}}$ and ${\mathcal{M}_{\rm L}}$.
@@
$\epsilon$=2.0 $\epsilon$=1.5 $\epsilon$=1.5 $\epsilon$=1.25 $\epsilon$=10.0 $\epsilon$=7.0
@@
D.4 Additional results of latent space analysis
Apart from the latent space discussion in Section 4.3, we show more analyses of other class triplets and datasets in this section.
Figure 17 presents an example of latent space visualization. Our primary observation is that, regardless of the base image’s class, ${\mathcal{D}_{\rm infer}}$ generated by LDPKiT-Sup is consistently more accurate compared to LDPKiT-Rand. As demonstrated in Figures 17(a) and 17(c), LDPKiT-Sup consistently forms a cluster whose distribution is closer to the target class distribution than LDPKiT-Rand. Furthermore, Figure 17(b) shows that LDPKiT-Sup’s cluster is positioned between the clusters of Class 4 and Class 8, while LDPKiT-Rand, despite being generated on Class 8 data, remains farther from the correct distribution. We validate these observations by calculating the Euclidean distances between cluster centroids. As shown in Table 9, while both LDPKiT-Sup and LDPKiT-Rand clusters are naturally closer to the Class 3 distribution, LDPKiT-Sup can generate a ${\mathcal{D}_{\rm infer}}$ cluster that aligns more closely with the target class of the base image. For instance, LDPKiT-Sup’s cluster is closest to Class 4 when superimposition is applied on two Class 4 data. In contrast, LDPKiT-Rand remains relatively static, always nearest to the Class 3 distribution regardless of the base image’s class. The results of SVHN in Figure 18 and Table 10 also align with other datasets. Visually, LDPKiT-Rand’s distribution remains relatively more static than LDPKiT-Sup, regardless of the label of the base image. LDPKiT-Sup, on the other hand, moves closer to the target cluster(s) of the two superimposed images.
Appendix E Model extraction attack detection with Zest
@@
E.1 Steps of model extraction detection
@@
    Calculate the Zest distance between the two models to compare, i.e., $D_{z}({\mathcal{M}_{\rm R}},{\mathcal{M}_{\rm L}})$, where ${\mathcal{M}_{\rm L}}$ is trained on the entire data split of the noisy ${\mathcal{D}_{\rm priv}}$, disjunctive to ${\mathcal{M}_{\rm R}}$’s training dataset.
@@
    Determine the existence of model extraction by comparing $\overline{D_{z}}$ with threshold 1.
    $\overline{D_{z}}<1$ indicates ${\mathcal{M}_{\rm R}}$ and ${\mathcal{M}_{\rm L}}$ are similar models and model extraction occurs.
    $\overline{D_{z}}>1$ indicates ${\mathcal{M}_{\rm R}}$ and ${\mathcal{M}_{\rm L}}$ are dissimilar, and thus no model extraction attack exists.
E.2 Zest distances supplementary results
We show the Cosine Distance measurement in the main paper as it is the most accurate metric [65]. We provide additional experimental results with $l_{1}$, $l_{2}$ and $l_{\infty}$ distance metrics supported by Zest in Tables 1112 and 13 on ${\mathcal{M}_{\rm L}}$ generated from Section 4.2. The results draw the same conclusion as Section 6.1 that the type of knowledge distillation in LDPKiT does not construct an adversarial model extraction attack quantitatively.

### v3

@@ A-C Hyperparameter choices and dataset preparation
A-D Additional experimental results
This section is complementary to the evaluation results presented in Section IV.
A-D1 Final accuracy on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$}
@@
Tables V, VI and VII show the tabulated final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of SVHN, Fashion-MNIST and SVHN that SIDP, LDPKiT-Rand and LDPKiT-Sup can achieve, which are also presented as bar graphs in Figure 4 in Section IV-B We record the accuracies at the last epoch of training. The results draw the same conclusion as Section IV that LDPKiT helps offset the accuracy trade-offs brought by LDP noise. The tables also show that LDPKiT can almost always achieve higher inference accuracy than SIDP, but the advantage is less obvious when the least noise is added (i.e., $\epsilon=30$). However, there is barely privacy protection in these cases, as SIDP accuracies are also very high, indicating a significant information leakage to $\mathcal{M}_{\rm R}$}. As more LDP noise is added to $\mathcal{D}_{\rm priv}$} to preserve privacy, the gap in utility that LDPKiT provides over SIDP also increases. Therefore, LDPKiT offers greater benefits in regimes with stronger privacy protection and correspondingly more noise.
A-D2 Additional Dataset Size Sensitivity Analysis
@@
The results in this section are complementary to the discussion in Section IV-D. Figure 11 demonstrates the impact of $|{\mathcal{D}_{\rm infer}}|$ with other $\epsilon$ values on the three image benchmarks. The takeaway aligns with the conclusion drawn from Figure 7 that although larger $\mathcal{D}_{\rm infer}$} helps model training and leads to higher accuracies, LDPKiT can already achieve satisfactory accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} when $|{\mathcal{D}_{\rm infer}}|$ is 500k. Table VIII shows the impact of $|{\mathcal{D}_{\rm infer}}|$ and $|{\mathcal{D}_{\rm priv}}|$ with other $\epsilon$ values. All the additional results align with the conclusions drawn from Table III. The main takeaway aligns with the discussion in Section IV-D. LDPKiT’s accuracies increase as $|{\mathcal{D}_{\rm infer}}|$ increases; however, the improvement becomes minor with larger $\mathcal{D}_{\rm infer}$} and the accuracies are sufficiently high when $|{\mathcal{D}_{\rm infer}}|$ is 500k. Furthermore, LDPKiT is less sensitive to $|{\mathcal{D}_{\rm priv}}|$, so LDPKiT can achieve decent accuracies when the user generates query data with a relatively smaller $\mathcal{D}_{\rm priv}$}.
A-D3 Qualitative Demonstration of Image Reconstruction
Figure 12 is complementary to the quantitative measures of image reconstruction success rate, which shows the reconstructed images from noisy data processed with LDPKiT-Sup when $\epsilon$ is low. The qualitative results also show that the reconstruction attack is not effective on data samples protected by LDPKiT-Sup. To demonstrate the authenticity of the denoiser, we also show the reconstruction results when $\epsilon$ is high (i.e., less noise added) in Figure 13.
@@ A-D3 Qualitative Demonstration of Image Reconstruction
TABLE IX: Euclidean distances between centroids of clusters shown in Figure 14 on PathMNIST.
@@
14(a) Sup (4,4) 1.9453 1.5981 2.5698
Rand 4 1.8724 2.5812 3.3226
14(b) Sup (8,4) 1.4984 2.1860 2.0151
Rand 8 1.6577 3.3427 3.0580
14(c) Sup (8,8) 1.6301 2.4562 1.5323
Rand 8 1.6407 3.3359 3.0413
@@
TABLE X: Euclidean distances between centroids of clusters shown in Figure 15 on SVHN.
@@
15(a) Sup (1,1) 2.6877 3.1031 3.2987
Rand 1 4.4212 3.0084 3.1381
15(b) Sup (1,8) 2.8854 3.0853 2.9836
Rand 1 4.4242 3.0094 3.1386
15(c) Sup (2,8) 3.8513 2.5689 3.1065
Rand 2 5.2321 3.0173 3.5754
@@
TABLE XI: Normalized Zest distance results with $l_{1}$ distance metric on $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$}.
@@
$\epsilon$=2.0
$\epsilon$=1.5
$\epsilon$=1.5
$\epsilon$=1.25
$\epsilon$=10.0
$\epsilon$=7.0
@@
TABLE XII: Normalized Zest distance results with $l_{2}$ distance metric on $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$}.
@@
$\epsilon$=2.0
$\epsilon$=1.5
$\epsilon$=1.5
$\epsilon$=1.25
$\epsilon$=10.0
$\epsilon$=7.0
@@
TABLE XIII: Normalized Zest distance results with $l_{\infty}$ distance metric on $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$}.
@@
$\epsilon$=2.0
$\epsilon$=1.5
$\epsilon$=1.5
$\epsilon$=1.25
$\epsilon$=10.0
$\epsilon$=7.0
@@
TABLE XIV: Estimated costs for image queries in USD. Google Vision uses per-request feature pricing; OpenAI and Azure assume low-detail vision images at 85 tokens/input image and \$5 per 1M input tokens.
# Image Queries Google Vision (\$0.0015/img) OpenAI Vision (85 tokens, \$0.000425/img) Azure OpenAI Vision (85 tokens, \$0.000425/img)
15,500 \$23.25 \$6.59 \$6.59
62,250 \$93.38 \$26.46 \$26.46
250,000 \$375.00 \$106.25 \$106.25
500,000 \$750.00 \$212.50 \$212.50
1,000,000 \$1,500.00 \$425.00 \$425.00
2,200,000 \$3,300.00 \$935.00 \$935.00
A-D4 Additional results of latent space analysis
Apart from the latent space discussion in Section IV-C, we show more analyses of other class triplets and datasets in this section.
Figure 14 presents an example of latent space visualization. Our primary observation is that, regardless of the base image’s class, $\mathcal{D}_{\rm infer}$} generated by LDPKiT-Sup is consistently more accurate compared to LDPKiT-Rand. As demonstrated in Figures 14(a) and 14(c), LDPKiT-Sup consistently forms a cluster whose distribution is closer to the target class distribution than LDPKiT-Rand. Furthermore, Figure 14(b) shows that LDPKiT-Sup’s cluster is positioned between the clusters of Class 4 and Class 8, while LDPKiT-Rand, despite being generated on Class 8 data, remains further from the correct distribution. We validate these observations by calculating the Euclidean distances between cluster centroids. As shown in Table IX, while both LDPKiT-Sup and LDPKiT-Rand clusters are naturally closer to the Class 3 distribution, LDPKiT-Sup can generate a $\mathcal{D}_{\rm infer}$} cluster that aligns more closely with the target class of the base image. For instance, LDPKiT-Sup’s cluster is closest to Class 4 when superimposition is applied to two Class 4 data points. In contrast, LDPKiT-Rand remains relatively static, always closest to the Class 3 distribution regardless of the base image’s class. The results of SVHN in Table X and Figure 15 align with our observations on other datasets. Visually, LDPKiT-Rand’s distribution remains relatively more static than LDPKiT-Sup, regardless of the label of the base image. LDPKiT-Sup, on the other hand, moves closer to the target cluster(s) of the two superimposed images.
A-E Model extraction attack detection with Zest
@@
A-E1 Steps of model extraction detection
@@
    Calculate the Zest distance between the two models to compare, i.e., $D_{z}({\mathcal{M}_{\rm R}},{\mathcal{M}_{\rm L}})$, where $\mathcal{M}_{\rm L}$} is trained on the entire data split of the noisy $\mathcal{D}_{\rm priv}$}, disjoint from $\mathcal{M}_{\rm R}$}’s training dataset.
@@
    Determine the existence of model extraction by comparing $\overline{D_{z}}$ with threshold 1.0 (as set in [99]). $\overline{D_{z}}<1.0$ indicates $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$} are similar models and model extraction exists.
    $\overline{D_{z}}>1.0$ indicates $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$} are dissimilar and model extraction attack does not exist.
A-E2 Zest distances supplementary results
We show the Cosine Distance measurement in the main paper as it is claimed to be the most accurate metric [99]. We provide additional experimental results with $l_{1}$, $l_{2}$ and $l_{\infty}$ distance metrics supported by Zest in Tables XIXII and XIII on $\mathcal{M}_{\rm L}$} generated from Section IV-B. None of the extraction performed in LDPKiT demonstrates a Zest distance <1.0<1.0. Therefore, the results draw the same conclusion as Section VI-A that the type of knowledge distillation in LDPKiT does not construct an adversarial model extraction attack quantitatively.
A-F Cost Analysis of LDPKiT on Major Service Providers
Based on the pricing information provided by Google [100], OpenAI [101] and Azure [102]. We estimate the cost of LDPKiT assuming the user makes $|{\mathcal{D}_{\rm infer}}|$ number of queries shown in Table III and Figure 7, ranging from 15,500 to 2.2M. The cost analysis is shown in Table XIV.